OpenStudy (anonymous):

find the final price of the item. cost:$12.50 Tax:7%

OpenStudy (saifoo.khan):

To find tax, \[12.50 + \frac{7}{100}\]

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

1.07(12.5)=$13.375

OpenStudy (anonymous):

13.375

OpenStudy (anonymous):

$12.50+12%($12.50) =1.07*$12.50 =$13.38

OpenStudy (anonymous):

I don't think you set the equation up correctly

OpenStudy (anonymous):

\[12.50(.07)+12.50= .088 +12.50 = 13.38\]